OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)?

OpenStudy (texaschic101):

y = mx + b slope(m) = 6 (-1,-8)...x = -1 and y = -8 now we sub -8 = 6(-1) + b -8 = -6 + b -8 + 6 = b -2 = b your equation is : y = 6x - 2

OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (1, 1) and has a slope of 2?

OpenStudy (texaschic101):

do it the same way.. y = mx + b slope(m) = 2 (1,1)..x = 1 and y = 1 now sub 1 = 2(1) + b 1 = 2 + b 1 - 2 = b -1 = b your equation is : y = 2x - 1

OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)? mc002-1.jpg mc002-2.jpg mc002-3.jpg mc002-4.jpg